OpenStudy (loser66):

verify that y_1 = e^x; y_2= cosh(x) and y_3= sinh(x) are solution of y"'-y"-y'+y=0 I got y_1. I got stuck at y_2 Please help

OpenStudy (loser66):

my work for y_2 is y= cosh(x) y'= -sinh(x) y" = -cosh(x) y"'= sinh(x) so, sinh(x)-(-cosh(x))-(-sinh(x)) +cosh(x)= 2sinh(x) +2cosh(x) \(\neq\)0

OpenStudy (loser66):

What's wrong with my stuff?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=d%2Fdx++cosh(x) How about this?

OpenStudy (loser66):

oops!! let me check on my note. hehe, just did by my bad memorize

OpenStudy (loser66):

Thank you very much,